i cannot figure this out!!!
" the second of two numbers is 1 more than twice the first. Their sum is 25. Find the numbers."
x=first number =
2x +1 = Second number =

Answer by Paul(988)

First number = x
Second number = 2x+1
Their sum = 25

(x)+(2x+1)=25
3x+1-1=25-1
3x=24
x=8 (first number)

2(8)+1=17 (second number)

Hence, the two numbers are 8 and 17.